# weighted-amenability and percolation

### v1

@@ Abstract.
The automorphism group of a transitive graph defines a weight function on the vertices through the Haar modulus. Benjamini, Lyons, Peres, and Schramm introduced the notion of weighted-amenability for a transitive graph, which is equivalent to the amenability of its automorphism group. We prove that this property is equivalent to level-amenability, that is, the property that the collection of vertices of weights in a given finite set always induces an amenable graph. We then use this to prove a version of Hutchcroft’s conjecture about $p_{h}<p_{u}$, relaxed à la Pak–Smirnova-Nagnibeda, where $p_{h}$ is the critical probability for the regime where clusters of infinite total weight arise, and $p_{u}$ is the uniqueness threshold. Further characterizations are given in terms of the spectral radius and invariant spanning forests. One consequence is the continuity of the phase transition at $p_{h}$ for weighted-nonamenable graphs.
@@ 1. Introduction and main results
Percolation theory began in 1957 by physicists [BH57], who were interested in studying the percolation of water in porous stones. Naturally, they considered a simplified version of the real-world phenomena by equating the stone with $\mathbb{Z}^{d}$ grid and letting the passages for water (i.e., bonds or edges) be open at random with probability $p$ independently of each other. Today it is a classical model and is known as Bernoulli$(p)$ percolation. By switching the focus from lattices and trees, in 1996 Benjamini and Schramm opened up the way to study percolation on general connected locally finite quasi-transitive graphs [BS96]. In turn, quasi-transitive graphs were further split into two categories: unimodular (those that “look like” Cayley graphs) and nonunimodular [HPS99, BLPS99b, Tim06b].
@@
For the class of nonunimodular transitive graphs, some standard techniques fail because of an intrinsic asymmetry. Thus, besides a few notable exceptions, many conjectures stated for all transitive graphs have been resolved only under the additional assumption of unimodularity. Transitive graphs are naturally equipped with a relative weight function, which is constant $1$ exactly for unimodular transitive graphs. Decomposition of vertices in sets of different weights quantifies the asymmetry in case of nonunimodular graphs. Besides posing a challenge from a technical perspective, another source of interest in this class of graphs is the particular phase transition phenomena that they produce [PPS06, Tim06b, Tan19]. Namely, due to the weights, the infinite subgraphs of nonunimodular graphs could be of two types: of infinite or finite total weight. We call such subgraphs heavy or light, respectively. Light infinite subgraphs often have properties somewhat in between finite and infinite components of the “usual” transitive graphs, thus exhibiting particularly interesting behavior. While often being more mysterious than their unimodular nonamenable counterparts, in some regards nonunimodular graphs are better understood. A breakthrough result by Hutchcroft [Hut20] shows the existence of several critical exponents for nonunimodular transitive graphs, and that they take mean-field values. This makes nonunimodular transitive graphs one of the few classes of transitive graphs where explicit critical exponents are known. It is important to note that even for unimodular graphs the existence of a nonunimodular subgroup of its automorphisms is useful, and in fact the mentioned results were proved in this generality. We will also follow the more general setup. Finally, another source of interest comes from Borel combinatorics and measured group theory, where an increasing amount of attention has been paid to the measure class preserving (mcp also known as quasi-pmp) Borel actions of countable groups [Tse22, TTD25, CTT22, BPZ24]. Establishing results in the mcp setting is in many ways parallel to studying percolation on nonunimodular transitive graphs because the orbit equivalence relation of such actions is also equipped with a relative weight function (known as the Radon–Nikodym cocycle). The connection between these fields goes beyond a similarity in themes and techniques, as results in one of them often enables the advances in the other.
@@
For the above reasons, we find it timely to revisit the notion of weighted-amenability of a graph and study its fundamental properties (see Definition 1.3). This notion was first introduced in [BLPS99b], where the authors showed its equivalence with the amenability of the automorphism group of the graph. For unimodular transitive graphs (where the weight function is constant $1$) it reduces to the classical notion of graph-amenability.
@@ 1.1. Percolation theory terminology
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$. For $x,y\in V$ that are in the same $\Gamma$-orbit we define the relative weight function $\mathbf{w}:V^{2}\to\mathbb{R}^{+}$ as
@@
where $\Gamma_{v}:=\{\gamma\in\Gamma\mid\gamma v=v\}$ is the stabilizer of $v\in V$. We say that the group $\Gamma$ acting on $G$ is unimodular if $\mathbf{w}^{y}(x)\equiv 1$, and the graph $G$ is unimodular if $\operatorname{Aut}(G)$ is. (For the standard definition of unimodularity via Haar measures and further details see Section 2.1.) We often refer to the weight of a vertex, fixing some arbitrary reference vertex $o$. Throughout this paper we omit writing the reference vertex $o$ in the superscript of $\mathbf{w}^{o}$, unless it is important to highlight. We also omit subscript $\Gamma$ when it is clear from the context. We say that a set of vertices is light (resp. heavy) if the sum of the weights is finite (resp. infinite). Notice that such definitions do not depend on the choice of the reference vertex. For unimodular graphs $\mathbf{w}\equiv 1$ any infinite cluster is automatically heavy. On the other hand, in the nonunimodular case, the light clusters could be either finite or infinite in size. Moreover, [HPS99, Theorem 4.1.6] implies that light-infinite clusters cannot coexist with heavy ones in Bernoulli percolation, yielding four percolation phases for any transitive graph (see Figure 1). Each of these phases can be nontrivial: see [HPS99, Tim06b, Tan19, Hut20] for details and examples.
@@
Since we are mostly concerned with the presence of weighs we do not restrict the scope of this paper to the weights induced by the full group of automorphisms. Instead, similarly to [BLPS99b, Hut20], we state our results for the weights induced by some closed $\Gamma\subseteq\operatorname{Aut}(G)$ that acts transitively on $G$. In particular, our analysis applies to unimodular transitive graphs whose automorphism group admits such a nonunimodular subgroup, with weights coming from the action of this subgroup.
@@
Figure 1. Depiction of four phases of Bernoulli$(p)$ percolation.
@@ 1.2. Weighted-amenability and its equivalents
[…]
@@
[…]
Notice that weighted-amenability does not depend on the choice of the reference vertex $o\in V$. It is also equivalent to taking infimum over all finite sets $F$, matching the original definition from [BLPS99b].
A relative weight function on a graph defines a decomposition of the vertex sets into levels. We say that two vertices belong to the same level if they are of equal weight. It turns out that weighted-amenability is equivalent to level-amenability, that is, the property that every finite union of levels induces an amenable subgraph. Level-amenability was first defined implicitly in [Tim06b], and it is at the center of Conjecture 8.5 of [Hut20].
@@
The equivalence of (1) and (3) was essentially shown in [BLPS99b]. More precise definitions and the proof are given in Section 4.
A further, probabilistic characterization of weighted-amenability will be Theorem 1.6, which generalizes the next result of Benjamini, Lyons, Peres, and Schramm.
@@
Following the above theorem, they asked if the converse to the last implication holds and showed that it does in the unimodular case. The second author answered their question negatively by constructing a counterexample in the nonunimodular case [Tim06b]. The original question of [BLPS99b] turns out to have a positive answer if we replace $p_{c}(\omega)$ with $p_{h}(\omega)$. Furthermore, the proper generalization of [BLPS99b, Theorem 5.3] is obtained by considering weights everywhere: $\mathbf{w}$-amenability replacing amenability, $p_{h}$ replacing $p_{c}$, and nonvanishing ends replacing ends. Then the respective parts of [BLPS99b, Theorem 5.3] (as in Theorem 1.5) turn into full equivalences:
@@
See Theorem 5.5 for some further equivalents, in the spirit of (2).
$\forall\alpha<1$, $\exists$ an invar. site perc. $\mathbf{P}$ with $\mathbf{P}(o\in\omega)>\alpha$ and no infinite cluster $\operatorname{Aut}(G)$ is amenable $G$ is $\mathbf{w}$-amenable $G$ is level-amenable $G$ is hyperfinite $p_{h}(G)=p_{u}(G)$exists increasing union of levels $\{G_{n}\}$ s.t. $\lim_{n\to\infty}p_{c}(G_{n})\geqslant p_{u}(G)$ The spectral radius of the $\sqrt{\mathbf{w}}$-biased RW $\rho^{(\mathbf{w})}(G)=1$exists an invariant random connected subgraph $\omega$ with $p_{h}(\omega)=1$exists invariant random spanning tree with $\leqslant$ 2 $\mathbf{w}$-nonvanishing ends All trees in any invariant random forest have $\leqslant$ 2 $\mathbf{w}$-nonvanishing ends All trees in any invariant random forest have $p_{h}=1$ Conjecture 6.2[BLPS99b, Theorem 3.9]Theorem 1.4[BLPS99b, Theorem 5.1][Hut20, Conjecture 8.5]Relaxation: Theorem 1.7[Tim06b, Corollary 5.8][Tan19, Corollary 4.8][BLPS99b, Theorem 5.1]Theorem 3.5 (Kesten)Theorem 1.6Theorem 5.5
Figure 2. Depiction of the relations between various characterisation of the weighted-amenability and relevant results.
@@ 1.3. Weighted-amenability and Bernoulli percolation
The fact that amenability of a graph implies $p_{c}=p_{u}$ was proven in [GKN92]; however the other direction of this conjecture is only known in particular cases that include transitive nonamenable graphs with suitably small spectral radius [BS96], large Cheeger constant [Sch01], with large girth [NP12], or for a suitably chosen Cayley graph of any nonamenable group [PSN00]. Finally, this conjecture was resolved for graphs whose automorphism group contains a nonunimodular subgroup acting quasi-transitively [Hut20]. In that work, Hutchcroft showed that for such graphs we have $p_{c}<p_{h}(\leqslant p_{u})$ and conjectured that level-amenability is equivalent to $p_{h}<p_{u}$ [Hut20, Conjecture 8.5]. While the forward direction was proven in [Tim06b] in 2006, Hutchcroft suggested that the other direction is likely harder to prove than the “$p_{c}<p_{u}$” conjecture of Benjamini and Schramm.
@@
In [Hut20], Hutchcroft also asked if it is true that if $\Gamma^{\prime}$ and $\Gamma\subseteq\Gamma^{\prime}\subseteq\operatorname{Aut}(G)$ are quasi-transitive subgroups of automorphisms for the locally finite connected graph $G$, then $p_{h}(G,\Gamma)>p_{h}(G,\Gamma^{\prime})$ if and only if $\mathbf{w}_{\Gamma}\neq\mathbf{w}_{\Gamma^{\prime}}$, [Hut20, Question 8.6]. We found a simple negative answer to this question, which shows that extra conditions are needed for the question. Consider an arbitrary transitive graph with a nonunimodular automorphism group, and add two extra leaves to each vertex, one yellow and one blue. Let $\Gamma^{\prime}$ be the set of all automorphisms of the new graph (where the colors of the leaves are ignored), and $\Gamma$ be all those automorphisms that respect the colors of the leaves. Then $\mathbf{w}_{\Gamma}\not=\mathbf{w}_{\Gamma^{\prime}}$, but their ratio is between $1/2$ and 2 everywhere, so $p_{h}(G,\Gamma)=p_{h}(G,\Gamma^{\prime})$. One way to ask the question now is to assume transitivity instead of quasi-transitivity. Another possibility is to require that the ratio of the two weight functions is unbounded. In the transitive case this is equivalent to them being distinct, for the following reason. Suppose that $G$ is transitive and $\mathbf{w}_{\Gamma}\not=\mathbf{w}_{\Gamma^{\prime}}$. We may assume that there is an edge $xy$ such that $\mathbf{w}_{\Gamma}(x)=\mathbf{w}_{\Gamma^{\prime}}(x)$ and $\mathbf{w}_{\Gamma}(y)\not=\mathbf{w}_{\Gamma^{\prime}}(y)$. Consider a $\gamma\in\Gamma$ that maps $x$ to $y$. Then the set $\{\mathbf{w}_{\Gamma}(\gamma^{n}(x))/\mathbf{w}_{\Gamma^{\prime}}(\gamma^{n}(x)):n\geq 1\}$ in unbounded.
@@
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $G$ is $\mathbf{w}$-nonamenable, then
@@
    there is no heavy cluster under Bernoulli$(p_{h}(G,\Gamma))$ bond percolation.

### v2

@@ Abstract.
In 1999, Benjamini, Lyons, Peres, and Schramm introduced a notion of weighted-amenability for transitive graphs that is equivalent to the amenability of its automorphism group. For unimodular graphs this notion coincides with classical graph-amenability and has been intensely studied. In the present work, we show that many classical unimodular results can be extended to the nonunimodular setting, which is further motivated by recent progress in the mcp (measure class preserving or quasi-pmp) setting of measured group theory. To this end, we prove new characterizations of weighted-amenability, in particular that it is equivalent to all finite unions of levels inducing amenable graphs. Hutchcroft conjectured that the latter property implies that $p_{h}<p_{u}$, where $p_{h}$ is the critical probability for the regime where clusters of Bernoulli percolation are infinite total weight and $p_{u}$ is the uniqueness threshold. We prove a relaxed version of his conjecture à la Pak–Smirnova-Nagnibeda. Further characterizations are given in terms of the spectral radius and invariant spanning forests. One of the consequences is the continuity of the phase transition at $p_{h}$ for weighted-nonamenable graphs.
@@ 1. Introduction and main results
Percolation theory began in 1957 by physicists [BH57], who were interested in studying the percolation of water in porous stones. Naturally, they considered a simplified version of the real-world phenomena by equating the stone with $\mathbb{Z}^{d}$ grid and letting the passages for water (i.e., bonds or edges) be open at random with probability $p$ independently of each other. Today it is a classical model and is known as Bernoulli(p)(p) percolation. By switching the focus from lattices and trees, in 1996 Benjamini and Schramm opened up the way to study percolation on general connected locally finite quasi-transitive graphs [BS96]. In turn, quasi-transitive graphs were further split into two categories: unimodular (those that “look like” Cayley graphs) and nonunimodular [HPS99, BLPS99b, Tim06b].
@@
For the class of nonunimodular transitive graphs, many standard techniques fail because of an intrinsic asymmetry. Thus, besides a few notable exceptions, many conjectures stated for all transitive graphs have been resolved only under the additional assumption of unimodularity. Transitive graphs are naturally equipped with a relative weight function, which is constant $1$ exactly for unimodular transitive graphs. Decomposition of vertices in sets of different weights quantifies the asymmetry in case of nonunimodular graphs. Besides posing a challenge from a technical perspective, another source of interest in this class of graphs is the particular phase transition phenomena that they produce [PPS06, Tim06b, Tan19, Hut20]. Namely, due to the weights, the infinite subgraphs of nonunimodular graphs could be of two types: where the total weight is infinite or finite. We call such subgraphs heavy or light, respectively. Light infinite subgraphs often have properties somewhat in between finite and infinite components of the “usual” transitive graphs, thus exhibiting particularly interesting behavior. While often being more mysterious than their unimodular nonamenable counterparts, in some regards nonunimodular graphs are better understood. A breakthrough result by Hutchcroft [Hut20] shows the existence of several critical exponents for nonunimodular transitive graphs, and that they take mean-field values. This makes nonunimodular transitive graphs one of the few classes of transitive graphs where explicit critical exponents are known. It is important to note that even for unimodular graphs the existence of a nonunimodular subgroup of its automorphisms is useful, and in fact the mentioned results were proved in this generality. We will also follow this more general setup. Finally, another source of interest comes from measured group theory and measurable combinatorics, where an increasing amount of attention has been paid to the measure class preserving (mcp also known as quasi-pmp) graphs, equivalence relations, and actions of countable groups [Tse22, TTD25, CTT22, BPZ24]. Establishing results in the mcp setting is in many ways parallel to studying percolation on nonunimodular transitive graphs through the relative weight function which is also defined for the orbit equivalence relation of such actions (and is known as the Radon–Nikodym cocycle). The connection between these fields extends beyond a similarity in themes and techniques, as results in one of them often enable advances in the other. The rising interest in the mcp setting is largely due to the success of the program initiated by Tserunyan in 2017 that aimed at filling the gaps in the understanding of such equivalence relations in comparison to their pmp counterpart. One of the highlights of this program is the recently announced resolution of a long standing question of measure equivalence classification of Baumslag-Solitar groups [GPT+25]. The present work develops the theory on the percolation side of the interplay and fits into the same program.
@@
For the above reasons, we find it timely to revisit the notion of weighted-amenability of a graph and study its fundamental properties (see Definition 1.3). This notion was first introduced in [BLPS99b], where the authors showed its equivalence to the amenability of the automorphism group of the graph. For unimodular transitive graphs (where the weight function is constant $1$) it reduces to the classical notion of graph-amenability.
@@ 1.1. Percolation theory terminology
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$. For $x,y\in V$ we define the relative weight function $\mathbf{w}:V^{2}\to\mathbb{R}^{+}$ as
@@
where $\Gamma_{v}:=\{\gamma\in\Gamma\mid\gamma v=v\}$ is the stabilizer of $v\in V$. We say that the group $\Gamma$ is unimodular if $\mathbf{w}^{y}(x)\equiv 1$, and the graph $G$ is unimodular if $\operatorname{Aut}(G)$ is. (For the standard definition of unimodularity via Haar measures and further details see Section 2.1.) We often refer to the weight of a vertex, fixing some arbitrary reference vertex $o$. Throughout this paper we omit writing the reference vertex $o$ in the superscript of $\mathbf{w}^{o}$, unless it is important to highlight. We also omit subscript $\Gamma$ when it is clear from the context. We say that a set of vertices is light (resp. heavy) if the sum of the weights is finite (resp. infinite). Notice that such definitions do not depend on the choice of the reference vertex. When $\Gamma$ is unimodular, i.e. $\mathbf{w}\equiv 1$, any infinite cluster is automatically heavy. On the other hand, in the nonunimodular case, the light clusters could be either finite or infinite in size. Moreover, [HPS99, Theorem 4.1.6] implies that light-infinite clusters cannot coexist with heavy ones in Bernoulli percolation, yielding four percolation phases for any transitive graph (see Figure 1). Each of these phases can be nontrivial: see [HPS99, Tim06b, Tan19, Hut20] for details and examples.
@@
Since we are mostly concerned with the presence of weighs we do not restrict the scope of this paper to the weights induced by the full group of automorphisms. Instead, similarly to [BLPS99b, Hut20], we state our results for the weights induced by some closed subgroup $\Gamma\subseteq\operatorname{Aut}(G)$ that acts transitively on $G$. In particular, our analysis applies to all transitive graphs, but yields new results only for those whose automorphism group admits such a nonunimodular subgroup.
@@
Figure 1. Depiction of four phases of Bernoulli(p)(p) percolation.
@@ 1.2. Weighted-amenability and its equivalents
where $\left|\cdot\right|$ denotes the size and $\partial_{V}$ denotes the external vertex boundary of a set.
@@
where $o$ is some reference vertex, $\left|\cdot\right|$ denotes the size, and $\partial_{V}$ denotes the external vertex boundary, and for a set $A\subseteq V$ we define $\mathbf{w}^{o}(A):=\sum_{x\in A}\mathbf{w}^{o}(x)$. Otherwise we call $G$ weighted-nonamenable or $\mathbf{w}$-nonamenable (with respect to the action of $\Gamma$).
Notice that weighted-amenability does not depend on the choice of the reference vertex $o\in V$. A similar notion of weighted isoperimetric constant was also introduced in the context of measured group theory by Kaimanovich [Kai97].
A relative weight function on a graph defines a decomposition of the vertex sets into levels. We say that two vertices belong to the same level if they are of equal weight. It turns out that weighted-amenability is equivalent to level-amenability, that is, the property that every finite union of levels induces an amenable subgraph. Level-amenability was first defined implicitly in [Tim06b], and it is at the center of [Hut20, Conjecture 8.5].
@@
[…]
Furthermore, we give a probabilistic characterization of weighted-amenability in Theorem 1.6, which generalizes the next result of Benjamini, Lyons, Peres, and Schramm.
@@
Following the above theorem, they asked if the converse to the last implication holds and showed that it does in the unimodular case. The second author answered their question negatively by constructing a counterexample in the nonunimodular case [Tim06b]. The original question of [BLPS99b] turns out to have a positive answer if we replace $p_{c}(\omega)$ with $p_{h}(\omega)$. Furthermore, the proper generalization of [BLPS99b, Theorem 5.3] is obtained by considering weights everywhere: $\mathbf{w}$-amenability replacing amenability, $p_{h}$ replacing $p_{c}$, and nonvanishing ends replacing ends. An end $\xi$ is called $\mathbf{w}$-vanishing if $\mathbf{w}(x_{n})$ converges to 0 for any sequence of vertices $x_{n}$ that converges to $\xi$. Then the respective parts of [BLPS99b, Theorem 5.3] (as in Theorem 1.5) turn into full equivalences:
@@
See Theorem 5.6 for some further equivalents, in the spirit of (2). In the proofs of Theorems 1.6 and 5.6 we rely on the connection with measured group theory (via the cluster graphing construction) and recent results in that field [TTD25, CTT22].
$\forall\alpha<1$, $\exists$ an invar. site perc. $\mathbf{P}$ with $\mathbf{P}(o\in\omega)>\alpha$ and no infinite cluster $\operatorname{Aut}(G)$ is amenable $G$ is $\mathbf{w}$-amenable $G$ is level-amenable $G$ is hyperfinite $p_{h}(G)=p_{u}(G)$ $\exists$ increasing union of levels $\{G_{n}\}$ s.t. $\lim_{n\to\infty}p_{c}(G_{n})\geqslant p_{u}(G)$ The spectral radius of the $\sqrt{\mathbf{w}}$-biased RW $\rho^{(\mathbf{w})}(G)=1$ $\exists$ an invariant random connected subgraph $\omega$ with $p_{h}(\omega)=1$ $\exists$ invariant random spanning tree with $\leqslant$ 2 $\mathbf{w}$-nonvanishing ends All trees in any invariant random forest have $\leqslant$ 2 $\mathbf{w}$-nonvanishing ends All trees in any invariant random forest have $p_{h}=1$ Conjecture 6.2[BLPS99b, Theorem 3.9]Theorem 1.4[BLPS99b, Theorem 5.1][Hut20, Conjecture 8.5]Relaxation: Theorem 1.7[Tim06b, Corollary 5.8][Tan19, Corollary 4.8][BLPS99b, Theorem 5.1]Theorem 3.5 (Kesten)Theorem 1.6Theorem 5.6
Figure 2. Depiction of the relations between various characterization of the weighted-amenability and relevant results.
@@ 1.3. Weighted-amenability and Bernoulli percolation
The fact that amenability of a graph implies $p_{c}=p_{u}$ was proven in [GKN92]; however the other direction of this conjecture is only known in particular cases that include transitive nonamenable graphs with suitably small spectral radius [BS96], large Cheeger constant [Sch01], with large girth [NP12], or for a suitably chosen Cayley graph of any nonamenable group [PSN00]. Finally, this conjecture was resolved for graphs whose automorphism group contains a nonunimodular subgroup acting quasi-transitively [Hut20]. In that work, Hutchcroft showed that for such graphs we have $p_{c}<p_{h}(\leqslant p_{u})$ and conjectured that level-amenability is equivalent to $p_{h}<p_{u}$ [Hut20, Conjecture 8.5]. While the forward direction was proven in [Tim06b] in 2006, Hutchcroft suggested that the other direction is likely harder to prove than the $p_{c}<p_{u}$” conjecture of Benjamini and Schramm.
@@
Let $\Gamma$ be a closed subgroup of $\operatorname{Aut}(G)$ that acts transitively on $G$ and $\mathbf{w}$ be the induced relative weight function as in (1). Suppose $G$ is $\mathbf{w}$-nonamenable. Then
@@
[…]
@@
In [Hut20], Hutchcroft also asked if it is true that if $\Gamma^{\prime}$ and $\Gamma\subseteq\Gamma^{\prime}\subseteq\operatorname{Aut}(G)$ are quasi-transitive subgroups of automorphisms of a locally finite connected graph $G$, then $p_{h}(G,\Gamma)>p_{h}(G,\Gamma^{\prime})$ if and only if $\mathbf{w}_{\Gamma}\neq\mathbf{w}_{\Gamma^{\prime}}$, [Hut20, Question 8.6]. We found a negative answer to this question via a simple example, which shows that extra conditions are needed for the question. Consider an arbitrary transitive graph with a nonunimodular automorphism group, and add two extra leaves to each vertex, one yellow and one blue. Let $\Gamma^{\prime}$ be the set of all automorphisms of the new graph (where the colors of the leaves are ignored), and $\Gamma$ be all those automorphisms that respect the colors of the leaves. Then $\mathbf{w}_{\Gamma}\not=\mathbf{w}_{\Gamma^{\prime}}$, but their ratio is between $1/2$ and 2 everywhere, so $p_{h}(G,\Gamma)=p_{h}(G,\Gamma^{\prime})$. One way to ask the question now is to assume transitivity instead of quasi-transitivity. Another possibility is to require that the ratio of the two weight functions is unbounded. In the transitive case this is equivalent to them being distinct, for the following reason. Suppose that $G$ is transitive and $\mathbf{w}_{\Gamma}\not=\mathbf{w}_{\Gamma^{\prime}}$. We may assume that there is an edge $xy$ such that $\mathbf{w}_{\Gamma}(x)=\mathbf{w}_{\Gamma^{\prime}}(x)$ and $\mathbf{w}_{\Gamma}(y)\not=\mathbf{w}_{\Gamma^{\prime}}(y)$. Consider a $\gamma\in\Gamma$ that maps $x$ to $y$. Then the set $\{\mathbf{w}_{\Gamma}(\gamma^{n}(x))/\mathbf{w}_{\Gamma^{\prime}}(\gamma^{n}(x)):n\geq 1\}$ in unbounded.